OpenStudy (anonymous):

the square of a sum of a number and 3 is 16. There are two solutions

OpenStudy (anonymous):

(x+3)^2 = 16 (x+3)(x+3) = 16 x^2 + 6x + 9 = 16 after subtracting 16 from each side you get... x^2 + 6x - 7 = 0 (x+7)(x-1) = 0 so either (x+7) = 0, or (x-1)=0 x+7 = 0 x+7-7 = 0-7 x = -7 or x-1 = 0 x-1+1 = 0+1 x = 1 so... x = -7 or x = 1

OpenStudy (dumbcow):

(x+3)^2 = 16 x+3 = +-4 x = 1 or -7